What length of tarpaulin 3 m wide will be required to make conical tent of height 8 m and base radius 6 m? Assume that the extra length of material that will be required for stitching margi,ns and wastage in cutting is approximately 20 cm. ( Use re = 3.14).

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Slant height of the tent = (6)2 + (8)2 m = 36 + 64 m = 10 m. Area of tarpaulin required = rrrl = 3.14 >< 6 x 10 m2 = 188.4 m2• .. Length or tarpaulin = 1s:•4 = 62.8 m. 190 Tut.al length required including wast.age = (62.8 + 0.2) m = 63 m.